The admission fee at a fair is $2 for children and $4 for adults. On a certain day, 2,000 people enter the fair and $6,000 is collected. How many children and how many adults went to see the fair on that day?
Just need some help! Thank you in advanced.

Answer by ankor@dixie-net.com(22740)

The admission fee at a fair is $2 for children and $4 for adults.
On a certain day, 2,000 people enter the fair and $6,000 is collected.
How many children and how many adults went to see the fair on that day?
:
Let a = no. of adults
Let c = no. of children
:
From the given information, we can derive two simple equations
the total no. of people
a + c = 2000
and
the total amt of $ received
4a + 2c = 6000
:
You can use substitution or elimination, we will use elimination
multiply the 1st equation by 2 and subtract from the 2nd equation
4a + 2c = 6000
2a + 2c = 4000
-----------------subtraction eliminates c, find a
2a = 2000
a = 2000/2
a = 1000 adults
:
It's a simple thing to find the no. of children, I'll let you do that
Check your solutions in the 2nd equation